# FlowAlign: Trajectory-Regularized, Inversion-Free Flow-based Image Editing

### v2

@@ 4 Experimental Results
Dataset, Baseline and Compute Resource. To verify the editing performance, we perform multiple analysis using PIE-Bench Ju et al. (2023) that contains 700 synthetic and natural images with paired original and editing prompts. For baseline algorithms, we focus on comparing with methods that establish trajectory between two samples. For the methods that utilizes noisy sample distribution to connect two trajectories, we use SDEdit Meng et al. (2021) and DDIB Su et al. (2023). For the method that improves the inversion process, we use RF-inversion Rout et al. (2025a). For the inversion-free method, we select FlowEdit Kulikov et al. (2024). Note that all these methods are training-free text-based image editing algorithms. For a fair comparison, we use the same flow model and 33 NFEs by following FlowEdit. In case of methods that involves inversion process, we use 17 NFEs for each inversion and sampling process. We conduct experiments using NVIDIA GeForce RTX 4090 (24GB VRAM). For runtime analysis, see Appendix D.
@@
Metric SDEdit Meng et al. (2021) DDIB Su et al. (2023) RF-Inv Rout et al. (2025a) FlowEdit Kulikov et al. (2024) Ours
@@
Human evaluation. Due to limitations of current metrics, such as CLIP similarity (as discussed in the previous section), we additionally conduct a human preference study. From a pool of 700 validation samples, we randomly select 100 images along with their corresponding original and editing text prompts. For each participant, we present a pairwise comparison between the edited image produced by the proposed method and that of a randomly selected baseline. Participants are asked to indicate which result they prefer, that is defined as one with accurate reflectance of editing prompt with source structure preservation. For a more detailed description of the human evaluation protocol, refer to Appendix C. Figure 3b summarizes the preference ratio of out method against each baseline. Across all comparisons, the proposed method is more preferred. These results demonstrates that our method achieves improved editing fidelity and superior source structure consistency, as intended.
@@
Backward Editing. In this work, we proposed a flow matching regularization to make smoother trajectory between two image samples. Thus, we investigate whether the proposed trajectory Eq. (26) behaves deterministically between two samples. To test this, we solve Eq. (26) in backward direction starting from the edited image. By evaluating the similarity between the source image and reconstructed image, we can evaluate whether the editing method indeed shows behavior like an ODE. We compute both pixel-wise metrics (PSNR, MSE) and perceptual metrics (LPIPS, DINO structural distance) and report the results in Table 1. Across all metrics, the proposed method outperforms the baselines. Theses results suggest that reverse editing using the proposed method nearly reconstructs the source image, supporting the effectiveness of the proposed trajectory regularization. Figure 5 shows a qualitative comparison for the edited image and reconstructed image. For example, in the last column, the proposed method uniquely reconstruct the torch of the statue to match the source image, whereas the baselines fail to reconstruct it accurately and instead convert it into a real torch. Importantly, the edited image obtained by the proposed method faithfully reflects the intended editing direction. This implies that the observed reconstruction ability stems from the smooth trajectory, rather than simply reducing changes during the editing process.
Ablation Study. The proposed method introduces two hyperparameters, $\omega$ and $\zeta$, which control the relative strength of each regularization term in Eq. (24). These parameters govern the trade-off between semantic alignment with the target prompt and structural consistency with the source image. To evaluate their impact, we conduct an ablation study under various settings. As shown in Figure 6, we observe that setting $\zeta=0.01$ consistently avhieves a favorable balance between semantic alignment and structure preservation. Notably, the framework without any additional gradients (i.e. $\omega=\zeta=0$, block dot) results in insufficient editing, which appears at the left-bottom position. This result emphasizes the effectiveness of the proposed CFG.
@@ 5 Conclusion
In this paper, we propose a flow matching regularization, that leads to smooth and stable editing trajectory, for inversion-free flow-based image editing algorithm. By defining differentiable optimal control problem with flow matching regularization, we explicitly balances semantic alignment with editing text and structural consistency with the source image. As a result, the proposed method achieves comparable or better editing performance while showing superior source consistency and better computational efficiency. Notably, we demonstrate that the simulated ODE exhibits a deterministic property by performing the reverse editing experiments. This work provides a novel design space for approximating ODEs between two samples without requiring additional training.
@@ Appendix A Optimal Control Formulation to Prove Proposition 1
Consider the following optimal control problem:
@@
$\ell({\bm{x}}_{t},{\bm{u}}_{t},t)$ $:=\frac{1}{2}\|{\bm{u}}_{t}-{\bm{a}}_{t}\|^{2}+\frac{\lambda}{2}\|{\bm{u}}_{t}-{\bm{b}}_{t}\|^{2}$ (28)
$m({\bm{x}}_{1})$ $:=\frac{\eta}{2}\|{\bm{x}}_{1}-{\bm{x}}_{end}\|^{2}$ (29)
@@
$H({\bm{x}}_{t},{\bm{u}}^{*}_{t},{\bm{p}}_{t},t):=$ $\frac{1}{2(1+\lambda)^{2}}\left(\|{\bm{p}}_{t}-\lambda({\bm{a}}_{t}-{\bm{b}}_{t})\|^{2}+\lambda\|{\bm{p}}_{t}-({\bm{b}}_{t}-{\bm{a}}_{t})\|^{2}\right)$
$+\frac{1}{1+\lambda}{\bm{p}}_{t}^{T}({\bm{a}}_{t}+\lambda{\bm{b}}_{t}-{\bm{p}}_{t})$ (34)
@@
$\dot{\bm{x}}_{t}$ $=\frac{\partial H({\bm{x}}_{t},{\bm{u}}^{*}_{t},{\bm{p}}_{t},t)}{\partial{\bm{p}}}$
$=\frac{1}{(1+\lambda)^{2}}({\bm{p}}_{t}-\lambda({\bm{a}}_{t}-{\bm{b}}_{t}))+\frac{\lambda}{(1+\lambda)^{2}}({\bm{p}}_{t}-({\bm{b}}_{t}-{\bm{a}}_{t}))+\frac{1}{1+\lambda}({\bm{a}}_{t}+\lambda{\bm{b}}_{t})-\frac{2}{1+\lambda}{\bm{p}}_{t}$
$=-\frac{1}{1+\lambda}{\bm{p}}_{t}+\frac{1}{1+\lambda}({\bm{a}}_{t}+\lambda{\bm{b}}_{t})$ (35)
@@
${\bm{x}}_{t_{0}}={\bm{x}}_{start},$ $\quad{\bm{p}}_{1}=\left.\frac{\partial m({\bm{x}}_{t},t)}{\partial{\bm{x}}_{t}}\right|_{t=1}=\eta({\bm{x}}_{1}-{\bm{x}}_{end})$ (37)
From Eq. (36), we can see that ${\bm{p}}_{t}$ is time-invariant constant, i.e. ${\bm{p}}_{t}={\bm{p}}$ for all $t\in[0,1]$. Combining this with Eq. (35), we have
${\bm{x}}_{1}$ $={\bm{x}}_{t_{0}}+\int_{t_{0}}^{1}-\frac{1}{(1+\lambda)}{\bm{p}}+\frac{1}{1+\lambda}({\bm{a}}_{t}+\lambda{\bm{b}}_{t})dt$ (38)
$=-\frac{1-t_{0}}{1+\lambda}{\bm{p}}+{\bm{x}}_{t_{0}}+{\bm{v}}_{t_{0}}$ (39)
@@
By plugging this in Eq. (37) with ${\bm{p}}_{1}={\bm{p}}$, we have
${\bm{p}}=\eta\left({\bm{x}}_{t_{0}}+{\bm{v}}_{t_{0}}-{\bm{x}}_{end}-\frac{(1-t_{0})}{(1+\lambda)}{\bm{p}}\right)$$\Rightarrow$ ${\bm{p}}=\frac{\eta}{1+\frac{\eta(1-t_{0})}{(1+\lambda)}}\left({\bm{x}}_{t_{0}}+{\bm{v}}_{t_{0}}-{\bm{x}}_{end}\right)$
@@
${\bm{u}}_{t}^{*}$ $=\frac{1}{1+\lambda}({\bm{a}}_{t}+\lambda{\bm{b}}_{t})-\frac{1}{1+\lambda}{\bm{p}}_{t}$ (41)
$=\frac{1}{1+\lambda}({\bm{a}}_{t}+\lambda{\bm{b}}_{t})-\frac{\eta}{(1+\lambda)+\eta(1-t)}\left({\bm{x}}_{t}+{\bm{v}}_{t}-{\bm{x}}_{end}\right)$ (42)
@@
${\bm{a}}_{t}:={\bm{v}}_{t}^{\theta}({\bm{p}}_{t},{\bm{c}}_{tgt})-{\bm{v}}_{t}^{\theta}({\bm{q}}_{t},{\bm{c}}_{src}),$ $\quad{\bm{b}}_{t}:={\bm{x}}_{tgt}-{\bm{x}}_{src},\quad\lambda\rightarrow\lambda_{t}$ (43)
and the time reversal $t\rightarrow 1-t$, leading to the modified boundary condition:
${\bm{x}}_{1}={\bm{x}}_{start}={\bm{x}}_{src},\quad{\bm{x}}_{0}={\bm{x}}_{end}={\bm{x}}_{tgt}$ (44)
This leads to the following:
@@
As $\eta\rightarrow 0$, this leads to the following ODE:
$\dot{\bm{x}}_{t}$ $=\frac{1}{1+\lambda_{t}}({\bm{a}}_{t}+\lambda_{t}{\bm{b}}_{t})={\bm{a}}_{t}+\frac{\lambda_{t}}{1+\lambda_{t}}({\bm{b}}_{t}-{\bm{a}}_{t})$ (46)
Furthermore, we have
${\bm{b}}_{t}-{\bm{a}}_{t}$ $={\bm{x}}_{tgt}-{\bm{x}}_{src}-\left({\bm{v}}_{t}^{\theta}({\bm{p}}_{t},{\bm{c}}_{tgt})-{\bm{v}}_{t}^{\theta}({\bm{q}}_{t},{\bm{c}}_{src})\right)$
$=\frac{(t{\bm{x}}_{tgt}+(1-t){\bm{x}}_{src})-{\bm{x}}_{src}}{t}-\left({\bm{v}}_{t}^{\theta}({\bm{p}}_{t},{\bm{c}}_{tgt})-{\bm{v}}_{t}^{\theta}({\bm{q}}_{t},{\bm{c}}_{src})\right)$
=(7)${{=}}\frac{{\bm{x}}_{t}-{\bm{x}}_{src}}{t}-\left({\bm{v}}_{t}^{\theta}({\bm{p}}_{t},{\bm{c}}_{tgt})-{\bm{v}}_{t}^{\theta}({\bm{q}}_{t},{\bm{c}}_{src})\right)$ )
=(11)${{=}}\frac{{\bm{p}}_{t}-{\bm{q}}_{t}}{t}-\left({\bm{v}}_{t}^{\theta}({\bm{p}}_{t},{\bm{c}}_{tgt})-{\bm{v}}_{t}^{\theta}({\bm{q}}_{t},{\bm{c}}_{src})\right)$ )
$=\frac{1}{t}\left(({\bm{p}}_{t}-t{\bm{v}}_{t}^{\theta}({\bm{p}}_{t},{\bm{c}}_{tgt})-({\bm{q}}_{t}-t{\bm{v}}_{t}^{\theta}({\bm{q}}_{t},{\bm{c}}_{src}))\right)$
$=\frac{1}{t}\left(\mathbb{E}[{\bm{p}}_{0}|{\bm{p}}_{t}]-\mathbb{E}[{\bm{q}}_{0}|{\bm{q}}_{t}]\right)$
where $\mathbb{E}[{\bm{p}}_{0}|{\bm{p}}_{t}]={\bm{p}}_{t}-t{\bm{v}}_{t}({\bm{p}}_{t},{\bm{c}}_{tgt})$ and $\mathbb{E}[{\bm{q}}_{0}|{\bm{q}}_{t}]={\bm{q}}_{t}-t{\bm{v}}_{t}({\bm{q}}_{t},{\bm{c}}_{src})$ refer to the Tweedie’s denoising estimates. Therefore, we have
$\dot{\bm{x}}_{t}={\bm{v}}_{t}({\bm{p}}_{t},{\bm{c}}_{tgt})-{\bm{v}}_{t}({\bm{q}}_{t},{\bm{c}}_{src})+\frac{\lambda_{t}}{(1+\lambda_{t})t}\left(\mathbb{E}[{\bm{p}}_{0}|{\bm{p}}_{t}]-\mathbb{E}[{\bm{q}}_{0}|{\bm{q}}_{t}]\right)$ (47)
@@ E.1 Quantitative and qualitative results
We provide the complete evaluation result on PIEBench, which includes metrics described in Section B. As discussed in the main paper, the proposed method outperforms all baselines in source consistency metrics while achieving a CLIP score comparable to FlowEdit. Although inversion-based editing algorithms such as DDIB and RF-Inversion exhibit higher CLIP scores, they tends to overemphasize the target concept at the expense of preserving the source structure (see more examples in Fig. 9 and 10) We also illustrate more qualitative results of the proposed method on text-based image editing task. Specifically, Fig. 9 and 10 present additional qualitative comparison between baselines and the FlowAlign. For diverse editing categories and objects, the proposed method shows better editing capability with better source structure preservation.
@@ F.2 Further Applications - Gaussian Splatting Editing
The main contribution of FlowAlign is to introducing flow matching regularization term for inversion-free image editing method, which has a goal of simulating the flow ODE between two image samples. Inspired by the proposed method, we also introduce a 3D editing via Gaussian splatting. The basic idea is to distill the knowledge of text-conditioned image prior of flow model to update Gaussians.
@@
$\nabla_{{\bm{x}}_{t}}\mathcal{L}_{FA}:={\bm{v}}_{t}({\bm{p}}_{t},{\bm{c}}_{tgt})-{\bm{v}}_{t}({\bm{q}}_{t},{\bm{c}}_{src})+\frac{\lambda_{t}}{(1+\lambda_{t})t}(\mathbb{E}[{\bm{p}}_{0}|{\bm{p}}_{t}]-\mathbb{E}[{\bm{q}}_{0}|{\bm{q}}_{t}]),$ (48)
@@
Because we use flow model defined in latent space, we compute the loss function using ${\bm{q}}_{t,c}=(1-t){\bm{x}}_{src,c}+t{\bm{\epsilon}}$, ${\bm{p}}_{t,c}={\bm{q}}_{t,c}+{\bm{x}}_{t,c}-{\bm{x}}_{src,c}$ where $c$ denotes sampled camera view, ${\bm{x}}_{src,c}=\mathcal{E}({\bm{g}}(\psi_{src},c))$ denotes the latent code of rendered view from initial Gaussian, ${\bm{x}}_{t,c}=\mathcal{E}({\bm{g}}(\psi,c))$ denotes the latent code of rendered vidw from current Gaussian, $\mathcal{E}$ denotes the pre-trained encoder of a VAE and ${\bm{\epsilon}}\sim\mathcal{N}(0,\bm{I})$. By using the chain rule, we obtain the following gradient for flow-based 3D editing:
@@
$=\left[{\bm{v}}_{t}({\bm{p}}_{t,c},{\bm{c}}_{tgt})-{\bm{v}}_{t}({\bm{q}}_{t,c},{\bm{c}}_{src})+\frac{\lambda_{t}}{(1+\lambda_{t})t}(\mathbb{E}[{\bm{p}}_{0}|{\bm{p}}_{t}]-\mathbb{E}[{\bm{q}}_{0}|{\bm{q}}_{t}])\right]\frac{\partial{\bm{x}}_{t}}{\partial\psi},$ (50)
@@
For the semantic guidance term in this gradient, we only apply CFG for ${\bm{p}}_{t}$ by setting ${\bm{c}}_{src}$ as null-text embedding, as same as the image editing algorithm.
We perform Gaussian Splatting optimization for $3,000$ iterations using a classifier-free guidance (CFG) weight $\omega$ set to either 70 or 100. We empirically set $\lambda_{t}=30\sigma_{t}^{2}$ where $\sigma_{t}=t$, which monotonically decreases toward zero as $t$ decreases from 0.9 to 0.1 as in the image editing case. Thus, flow-matching guidance is emphasized during the early Gaussian updates. We set the step size $\eta_{t}=1+30\sigma_{t}^{2}$. This setting consistently achieves a favorable balance between semantic fidelity and structural integrity. Experiments are conducted on real-world scenes using datasets from IN2N [7] and PDS [15]. For comparison, we benchmark against existing distillation methods, including SDS [26], DDS [8], and PDS [15], as well as the non-distillation baseline IGS2GS [34]. For baselines using distillation methods we used Stable Diffusion 3.0 and for IGS2GS [34], we use UltraEdit [36] in replacement of Instruct-Nerf2Nerf  [7] for flow-based models. One can find the pseudocode for the 3D editing with FlowAlign in Algorithm 2, and additional results on figure15 and figure 15.
@@
1:Source parameter $\psi_{src}$, Pre-trained flow-based model ${\bm{v}}_{\theta}$, VAE encoder $\mathcal{E}$, Source/Target text embeddings ${\bm{c}}_{src},{\bm{c}}_{tgt}$, CFG scale $\omega$, camera views $Cams$, Noise Schedule $\sigma_{t}$, differentiable generator ${\bm{g}}$, min and max timestep $T_{min},T_{max}$, time-varying coefficient $\lambda_{t}$, step size $\eta_{t}$
@@
12:     $\nabla_{\psi}\mathcal{L}_{FA}\leftarrow[({\bm{v}}_{\theta}({\bm{p}}_{t,c})-{\bm{v}}_{\theta}({\bm{q}}_{t,c})+\frac{\lambda_{t}}{(1+\lambda_{t})t}(\mathbb{E}[{\bm{p}}_{0}|{\bm{p}}_{t}]-\mathbb{E}[{\bm{q}}_{0}|{\bm{q}}_{t})){\frac{\partial{\bm{x}}}{\partial\psi}}]$

### v3

@@ 4 Experimental Results
Dataset, Baseline and Compute Resource. To verify the editing performance, we perform multiple analysis using PIE-Bench DBLP:journals/corr/abs-2310-01506 that contains 700 synthetic and natural images with paired original and editing prompts. For baseline algorithms, we focus on comparing with methods that establish trajectory between two samples. For the methods that utilizes noisy sample distribution to connect two trajectories, we use SDEdit meng2021sdedit and DDIB su2022dual . For the method that improves the inversion process, we use RF-inversion rout2025semantic . For the inversion-free method, we select FlowEdit kulikov2024flowedit . Note that all these methods are training-free text-based image editing algorithms. For a fair comparison, we use the same flow model and 33 NFEs by following FlowEdit. In case of methods that involves inversion process, we use 17 NFEs for each inversion and sampling process. We conduct experiments using NVIDIA GeForce RTX 4090 (24GB VRAM). For runtime analysis, see Appendix D.
@@
Metric SDEdit meng2021sdedit DDIB su2022dual RF-Inv rout2025semantic FlowEdit kulikov2024flowedit Ours
@@
Human evaluation. Due to limitations of current metrics, such as CLIP similarity (as discussed in the previous section), we additionally conduct a human preference study.
From a pool of 700 validation samples, we randomly select 100 images along with their corresponding original and editing text prompts. For each participant, we present a pairwise comparison between the edited image produced by the proposed method and that of a randomly selected baseline. Participants are asked to indicate which result they prefer, that is defined as one with accurate reflectance of editing prompt with source structure preservation. For a more detailed description of the human evaluation protocol, refer to Appendix C. Figure 3b summarizes the preference ratio of out method against each baseline. Across all comparisons, the proposed method is more preferred. These results demonstrates that our method achieves improved editing fidelity and superior source structure consistency, as intended.
Backward Editing. In this work, we proposed a flow matching regularization to make smoother trajectory between two image samples. Thus, we investigate whether the proposed trajectory Eq. (24) behaves deterministically between two samples. To test this, we solve Eq. (24) in backward direction starting from the edited image. By evaluating the similarity between the source image and reconstructed image, we can evaluate whether the editing method indeed shows behavior like an ODE. We compute both pixel-wise metrics (PSNR, MSE) and perceptual metrics (LPIPS, DINO structural distance) and report the results in Table 1. Across all metrics, the proposed method outperforms the baselines. These results suggest that reverse editing using the proposed method nearly reconstructs the source image, supporting the effectiveness of the proposed trajectory regularization. Figure 5 shows a qualitative comparison for the edited image and reconstructed image. For example, in the last column, the proposed method uniquely reconstruct the torch of the statue to match the source image, whereas the baselines fail to reconstruct it accurately and instead convert it into a real torch. Importantly, the edited image obtained by the proposed method faithfully reflects the intended editing direction. This implies that the observed reconstruction ability stems from the smooth trajectory, rather than simply reducing changes during the editing process.
@@
Ablation Study. The proposed method introduces two hyperparameters, $\omega$ and $\zeta$, which control the relative strength of each regularization term in Eq. (22). These parameters govern the trade-off between semantic alignment with the target prompt and structural consistency with the source image. To evaluate their impact, we conduct an ablation study under various settings. As shown in Figure 6, we observe that setting $\zeta=0.01$ consistently achieves a favorable balance between semantic alignment and structure preservation. Notably, the framework without any additional gradients (i.e. $\omega=\zeta=0$, block dot) results in insufficient editing, which appears at the left-bottom position. This result emphasizes the effectiveness of the proposed CFG.
@@ 5 Conclusion
In this paper, we propose a flow matching regularization, that leads to smooth and stable editing trajectory, for inversion-free flow-based image editing algorithm. By defining differentiable optimal control problem with similarity regularization at the terminal point, we explicitly balances semantic alignment with editing text and structural consistency with the source image. As a result, the proposed method achieves comparable or better editing performance while showing superior source consistency and better computational efficiency. Notably, we demonstrate that the simulated ODE exhibits a deterministic property by performing the reverse editing experiments. In summary, this work provides a novel design space for approximating ODEs between two samples without requiring additional training.
@@ Appendix A Optimal Control Formulation to Prove Proposition 1
Consider the following time-reversal optimal control problem:
@@
$\ell({\bm{x}}_{t},{\bm{u}}_{t},t)$ $:=\frac{1}{2}\|{\bm{u}}_{t}-{\bm{a}}_{t}\|^{2}$ (26)
$m({\bm{x}}_{1})$ $:=\frac{\eta}{2}\|{\bm{x}}_{0}-{\bm{x}}_{src}\|^{2}$ (27)
@@
$H({\bm{x}}_{t},{\bm{u}}^{*}_{t},{\bm{p}}_{t},t):=$ $-\frac{1}{2}\|{\bm{p}}_{t}\|^{2}+{\bm{p}}_{t}^{T}{\bm{a}}_{t}$ (32)
@@
$\dot{\bm{x}}_{t}$ $={\bm{u}}_{t}^{*}={\bm{a}}_{t}-{\bm{p}}_{t}$ (33)
@@
${\bm{x}}_{t_{0}}={\bm{x}}_{start},$ $\quad{\bm{p}}_{0}=\left.\frac{\partial m({\bm{x}}_{t},t)}{\partial{\bm{x}}_{t}}\right|_{t=0}=\eta({\bm{x}}_{0}-{\bm{x}}_{src})$ (35)
From Eq. (34), we can see that ${\bm{p}}_{t}$ is time-invariant constant, i.e. ${\bm{p}}_{t}={\bm{p}}$ for all $t\in[0,1]$. Accordingly, we have
${\bm{x}}_{0}$ $={\bm{x}}_{t_{0}}+\int_{t_{0}}^{0}-{\bm{p}}+{\bm{a}}_{t}dt$ (36)
$=t_{0}{\bm{p}}+{\bm{x}}_{t_{0}}+{\bm{v}}_{t_{0}}$ (37)
@@
By plugging this in Eq. (35) with ${\bm{p}}_{0}={\bm{p}}$, we have
${\bm{p}}=\eta\left(t_{0}{\bm{p}}+{\bm{x}}_{t_{0}}+{\bm{v}}_{t_{0}}-{\bm{x}}_{src}\right)$ $\quad\Rightarrow{\bm{p}}=\frac{\eta({\bm{x}}_{t_{0}}+{\bm{v}}_{t_{0}}-{\bm{x}}_{src})}{1-t_{0}\eta}$
@@
${\bm{u}}_{t}^{*}$ $={\bm{a}}_{t}+\frac{-\eta}{1-\eta t}\left({\bm{x}}_{t}+{\bm{v}}_{t}-{\bm{x}}_{src}\right)$ (39)
@@
${\bm{a}}_{t}:={\bm{v}}_{t}^{\theta}({\bm{p}}_{t},{\bm{c}}_{tgt})-{\bm{v}}_{t}^{\theta}({\bm{q}}_{t},{\bm{c}}_{src})$ (40)
This leads to the following first order approximation:
@@
$\simeq-t{\bm{v}}_{t}({\bm{p}}_{t})+t{\bm{v}}_{t}({\bm{q}}_{t})$ (42)
Accordingly, we have
${\bm{x}}_{t}+{\bm{v}}_{t}-{\bm{x}}_{src}$ $={\bm{p}}_{t}-{\bm{q}}_{t}+{\bm{v}}_{t}$
$\simeq{\bm{p}}_{t}-{\bm{q}}_{t}+(-t{\bm{v}}_{t}({\bm{p}}_{t})+t{\bm{v}}_{t}({\bm{q}}_{t}))$
$=\mathbb{E}[{\bm{p}}_{0}|{\bm{p}}_{t}]-\mathbb{E}[{\bm{q}}_{0}|{\bm{q}}_{t}]$
where we use ${\bm{x}}_{t}-{\bm{x}}_{src}={\bm{p}}_{t}-{\bm{q}}_{t}$ from Eq. (11) for the first equality. Thus, the optimal control ${\bm{u}}_{t}^{*}$ is given by
${\bm{u}}_{t}^{*}$ $={\bm{a}}_{t}+\frac{-\eta}{1-\eta t}\left({\bm{x}}_{t}+{\bm{v}}_{t}-{\bm{x}}_{src}\right)$
$\simeq{\bm{a}}_{t}+\frac{-\eta}{1-\eta t}\left(\mathbb{E}[{\bm{p}}_{0}|{\bm{p}}_{t}]-\mathbb{E}[{\bm{q}}_{0}|{\bm{q}}_{t}]\right)$ (43)
Therefore, we have
$\dot{\bm{x}}_{t}$ $={\bm{v}}_{t}({\bm{p}}_{t},{\bm{c}}_{tgt})-{\bm{v}}_{t}({\bm{q}}_{t},{\bm{c}}_{src})+\frac{-\eta}{1-\eta t}\left(\mathbb{E}[{\bm{p}}_{0}|{\bm{p}}_{t}]-\mathbb{E}[{\bm{q}}_{0}|{\bm{q}}_{t}]\right)$ (44)
$={\bm{v}}_{t}({\bm{p}}_{t},{\bm{c}}_{tgt})-{\bm{v}}_{t}({\bm{q}}_{t},{\bm{c}}_{src})+\gamma\left(\mathbb{E}[{\bm{p}}_{0}|{\bm{p}}_{t}]-\mathbb{E}[{\bm{q}}_{0}|{\bm{q}}_{t}]\right)$ (45)
@@ E.1 Quantitative and qualitative results
We provide the complete evaluation result on PIEBench, which includes metrics described in Section B. As discussed in the main paper, the proposed method outperforms all baselines in source consistency metrics while achieving a CLIP score comparable to FlowEdit. Although inversion-based editing algorithms such as DDIB and RF-Inversion exhibit higher CLIP scores, they tends to overemphasize the target concept at the expense of preserving the source structure (see more examples in Fig. 9 and 10). We also illustrate more qualitative results of the proposed method on text-based image editing task. Specifically, Fig. 9 and 10 present additional qualitative comparison between baselines and the FlowAlign. For diverse editing categories and objects, the proposed method shows better editing capability with better source structure preservation.
@@ F.2 Further Applications - Gaussian Splatting Editing
Recall that one of the main contributions of FlowAlign is introducing a regularization term for inversion-free image editing method, which has a goal of simulating the flow ODE between two image samples. Accordingly, we can apply FlowAlign for a 3D editing via Gaussian splatting. The basic idea is to distill the knowledge of text-conditioned image prior of flow model to update Gaussians.
@@
$\nabla_{{\bm{x}}_{t}}\mathcal{L}_{FA}:={\bm{v}}_{t}({\bm{p}}_{t},{\bm{c}}_{tgt})-{\bm{v}}_{t}({\bm{q}}_{t},{\bm{c}}_{src})+\gamma(\mathbb{E}[{\bm{p}}_{0}|{\bm{p}}_{t}]-\mathbb{E}[{\bm{q}}_{0}|{\bm{q}}_{t}]),$ (46)
@@
Because we use flow model defined in latent space, we compute the loss function using ${\bm{q}}_{t,c}=(1-t){\bm{x}}_{src,c}+t{\bm{\epsilon}}$, ${\bm{p}}_{t,c}={\bm{q}}_{t,c}+{\bm{x}}_{t,c}-{\bm{x}}_{src,c}$. Here, $c$ denotes sampled camera view, ${\bm{x}}_{src,c}=\mathcal{E}({\bm{g}}(\psi_{src},c))$ denotes the latent code of rendered view from initial Gaussian, ${\bm{x}}_{t,c}=\mathcal{E}({\bm{g}}(\psi,c))$ denotes the latent code of rendered vidw from current Gaussian, $\mathcal{E}$ denotes the pre-trained encoder of a VAE and ${\bm{\epsilon}}\sim\mathcal{N}(0,\bm{I})$. By using the chain rule, we obtain the following gradient for flow-based 3D editing:
@@
$=\left[{\bm{v}}_{t}({\bm{p}}_{t,c},{\bm{c}}_{tgt})-{\bm{v}}_{t}({\bm{q}}_{t,c},{\bm{c}}_{src})+\gamma(\mathbb{E}[{\bm{p}}_{0}|{\bm{p}}_{t}]-\mathbb{E}[{\bm{q}}_{0}|{\bm{q}}_{t}])\right]\frac{\partial{\bm{x}}_{t}}{\partial\psi},$ (48)
@@
For the semantic guidance term in this gradient, we only apply CFG for ${\bm{p}}_{t}$ by setting $c_{src}$ as null-text embedding, as same as the image editing algorithm.
We perform Gaussian Splatting optimization for $3,000$ iterations using a classifier-free guidance (CFG) weight $\omega$ set to either 70 or 100. Experiments are conducted on real-world scenes using datasets from IN2N [7] and PDS [15]. For comparison, we benchmark against existing distillation methods, including SDS [26], DDS [8], and PDS [15], as well as the non-distillation baseline IGS2GS [34]. For baselines using distillation methods we used Stable Diffusion 3.0 and for IGS2GS [34], we use UltraEdit [36] in replacement of Instruct-Nerf2Nerf  [7] for flow-based models. One can find the pseudocode for the 3D editing with FlowAlign in Algorithm 2, and additional results on Figure 15 and Figure 15.
@@
1:Source parameter $\psi_{src}$, Pre-trained flow-based model ${\bm{v}}_{\theta}$, VAE encoder $\mathcal{E}$, Source/Target text embeddings ${\bm{c}}_{src},{\bm{c}}_{tgt}$, CFG scale $\omega$, camera views $Cams$, Noise Schedule $\sigma_{t}$, differentiable generator ${\bm{g}}$, min and max timestep $T_{min},T_{max}$
@@
12:     $\nabla_{\psi}\mathcal{L}_{FA}\leftarrow[({\bm{v}}_{\theta}({\bm{p}}_{t,c})-{\bm{v}}_{\theta}({\bm{q}}_{t,c})+\gamma(\mathbb{E}[{\bm{p}}_{0}|{\bm{p}}_{t}]-\mathbb{E}[{\bm{q}}_{0}|{\bm{q}}_{t})){\frac{\partial{\bm{x}}}{\partial\psi}}]$